OpenStudy (anonymous):

differentiate d/dt (a + (2bt/(1-bt^2))) need help with the rules.. :/

OpenStudy (anonymous):

\[\frac{ d }{ dt } [a+\frac{ 2bt }{ (1-bt ^{2} )}]\]

OpenStudy (irishboy123):

\(a,b = const \implies (a)' = 0\) so you're off to a flying start....

OpenStudy (anonymous):

ya.. and i now the (1-bt^2)^2 will be in the denominator, * -2bt, but I don't know what to do with the numerator

OpenStudy (irishboy123):

for \(\large \frac{ d }{ dt } [\frac{ 2bt }{ (1-bt ^{2} )}]\), most people would use the quotient rule \(\left( \dfrac{u}{v}\right)^{\prime} = \dfrac{u'v - uv'}{v^2}\)

OpenStudy (anonymous):

ahhhh okay... thats what i was looking for. i need a book with all these dang cal rules.

OpenStudy (irishboy123):

or you can re-write as \(\large \frac{ d }{ dt } [ 2bt (1-bt ^{2} )^{-1}]\) and use product rule....

OpenStudy (anonymous):

can you give me the product rule by chance so i can write it down.

OpenStudy (irishboy123):

\((uv)' = u'v + uv'\)

OpenStudy (irishboy123):

here's a very comprehensive cheat sheet from a very well regarded source http://tutorial.math.lamar.edu/pdf/Calculus_Cheat_Sheet_All.pdf

OpenStudy (anonymous):

Thank you, appreciate it. I know I should remember all this stuff already, which is why it probably isn't in the book.

OpenStudy (irishboy123):

good luck:-)

OpenStudy (anonymous):

do you see my error?? @IrishBoy123

OpenStudy (johnweldon1993):

Ahh you just missed a negative sign on the bottom \[\large 2bt(1 - bt^2)^{-1}\] \[\large u'v + uv' = 2b(1 - bt^2)^{-1} + 2bt(\color \red{-}(1-bt^2)^{-2})(-2bt))\]

OpenStudy (irishboy123):

sorry zona and thanks john i was occupied over on physics....

OpenStudy (johnweldon1993):

Yeah no problem, easy to get occupied over there...physics is that great after all :D

OpenStudy (anonymous):

ohhh right bring the -1 in front then subtract 1 from the exponent... silly mistake. thank you guys.